Each of the 30 students in homeroom 704 contributed either a nickel or a quarter to the Cancer Fund. If the total amount collected was $4.70. how many students contributed a nickel?
Answer by macston(5194)

.
N=number of nickels; Q=number of quarters=30-N
.
$0.05N+$0.25Q=$4.70
$0.05N+$0.25(30-N)=$4.70
$0.05N+$7.50-$0.25N=$4.70
-$0.20N=-$2.80
N=14
.
ANSWER: 14 students contributed a nickel.
.
CHECK:
If 14 gave a nickel, 30-14=16 gave a quarter.
$0.05N+$0.25Q=$4.70
$0.05(14)+$0.25(16)=$4.70
$0.70+$4.00=$4.70
$4.70=$4.70
.